Linda must choose a number between 61 and 107 that is a multiple of 2,3, and 9. Write all the numbers that she should choose if there is more than one number separate them with commas.
can anyone help? thanks, don't have to show work:D

Answer by ewatrrr(24785)


Hi
The number being a multiple of 9 is KEY, as well as being a multiple of '2'
9*8 = 72, 9*10 = 90 (63,81,99 will not work)
72, 90